OpenStudy (anonymous):

what is the corresponding integration formula : (3x^2)/(2*(x^3+5)^(1/2))

OpenStudy (anonymous):

Solution: \[\int\frac{3x^2}{\sqrt{2x^3+5}}dx=\sqrt{2x^3+5}+const.\]

OpenStudy (anonymous):

let u= 2x^3+5, and solve it by u substitution